i need help on solving equations using addition and attraction. example -2=x-4.
Answer by pushpaharan(47)

In solving equation, first simplify both sides as mush as you can. then move around the numbers and variables such that all variables on one side and numbers on the other side. follow the example.
-2=x-4
-2+4 = x
2=x
therefore x=2
follow these rules
when a positive number is moved to other side of the equation it will become negative number
x+5 =7
x= 7-5
when a negative number is moved to other side of the equation it will become positive number
x-3 = -4
x= -4 +3
x=-1
when a multiplied number is moved to other side of the equation it will become quotient (divisor)
2*x=10
x=10/2
x=5
when a divided number (quotient) is moved to other side of the equation it will become multiplied number.
x/4 =5
x= 5*4
x=20
also make sure in complex question you may need to follow the order of operation (BEDMAS)
What a long answer :-)